What is the relationship between the number of vertices, the number of faces and the number of edges of a triangular prism and a pyramid whose base is a regular triangle?
Answer by ikleyn(52794)

.
What is the relationship between the number of vertices, the number of faces and the number of edges
of a triangular prism and a pyramid whose base is a regular triangle?
---------------------------------------------------------------

For a triangular prism: F - E + V = 5 - 9 + 6 = 2,
    where F is the number of faces, E is the number of edges and V is the number of vertices.

For a triangular pyramid: F - E + V = 4 - 6 + 4 = 2,
    the same number as for a triangular prism.

It is the consequence of the famous  Euler's theorem.

The fact that the base is a regular triangle,  is not essentially important.
The relation is valid for many polyhedra that are surfaces of convex solid bodies.  See the referred Wikipedia article.